An electron,proton and alpha particle have same linear momenta . Which one of them posseses least kinetic energy?

Grade:12

1 Answers

poonam
6 years ago
K.E =1/2 mv2P=mv so KE = p^2/2m. As the linear momentum is same so KE will depend on their mass inversely. So as alpha particle has highest mass among these three so it will have least KE